# Learning functions on symmetric matrices and point clouds via lightweight invariant features

### v2

@@ Abstract
In this work, we present a mathematical formulation for machine learning of (1) functions on symmetric matrices that are invariant with respect to the action of permutations by conjugation, and (2) functions on point clouds that are invariant with respect to rotations, reflections, and permutations of the points. To achieve this, we construct $O(n^{2})$ invariant features derived from generators for the field of rational functions on $n\times n$ symmetric matrices that are invariant under joint permutations of rows and columns. We show that these invariant features can separate all distinct orbits of symmetric matrices except for a measure zero set; such features can be used to universally approximate invariant functions on almost all weighted graphs. For point clouds in a fixed dimension, we prove that the number of invariant features can be reduced, generically without losing expressivity, to $O(n)$, where $n$ is the number of points. We combine these invariant features with DeepSets to learn functions on symmetric matrices and point clouds with varying sizes. We empirically demonstrate the feasibility of our approach on molecule property regression and point cloud distance prediction.
@@ 1 Introduction
Motivated by such considerations of computational scalability, we propose a relaxation of universal approximation. At the cost of throwing out a closed, measure zero set of “bad” point clouds (respectively, weighted graphs), we provide an approach to these learning problems that has a universal approximation guarantee on the remaining “good” point clouds (respectively, weighted graphs), and has computational cost bounded by a constant multiple of the size of the input data.
Our method is to look for computationally tractable invariants that contain the same information as a set of generators for the field of invariant rational functions on the data. It is known that such field generators separate orbits away from a Zariski-closed (and thus Euclidean-closed and measure zero) “bad” subset; thus the proposed invariants do as well. A standard argument based on the Stone-Weierstrass theorem then implies that the proposed invariants are universally approximating away from the “bad” subset.
@@
  • Using Galois theory, we provide $O(n^{2})$ invariant features that generically separate real symmetric matrices with respect to the action (1), so can be used to universally approximate invariant functions on almost all such matrices (Theorem 3.1).
  • We show that the same features, applied to the Gram matrix of a point cloud, also generically separate (and therefore generically universally approximate invariant functions on) point clouds $V\in\mathbb{R}^{d\times n}$ with respect to the natural $\mathrm{O}(d)\times\mathfrak{S}_{n}$ action (Theorem 4.1).
  • To improve efficiency, we prove that in the point cloud setting we can express an equivalent class of functions using $O(n)$ invariant features for fixed $d$ (Theorem 5.5).
@@ 2 Related work
Invariant theory. Describing the class of functions invariant under a group action is historically the project of invariant theory. Other works to consider, from the point of view of invariant theory, the same or closely related group actions to the ones discussed above, include [ACG96, Thi00, KLR22, GHL21, Had21]; the last two of these have applications to ML in mind. These papers compute complete sets of invariants (i.e., generators for the algebra of invariant polynomials), or separating sets (i.e., invariant polynomials whose evaluation can classify orbits to the extent possible) in special cases. However, a complete or even a separating set of invariants for general $n,d$ is expected to be computationally intractable, because such a set would solve the graph isomorphism problem, for which the fastest known algorithm has quasi-polynomial time complexity. The starting point for the present investigation was the recognition that the method of [Thi00, Theorem 11.2]—which gives generators for the field of rational invariants in a closely related situation where the algebra generators remain out of reach—could be used to produce a set of computationally tractable invariants that uniquely identifies a generic (though not a worst-case) symmetric matrix, up to the $\mathfrak{S}_{n}$ conjugation action, and is not meaningfully bigger than the size of the input data. Other works that use field generation as a relaxation of algebra generation to lower a computational hurdle include [BBSK+23, BSGHR24]. More general work on computation of fields of rational invariants includes [MB99, Kem07, HK07, HL12, KK12, HL16, Kem16].
@@
Machine learning on weighted graphs. Graph neural networks (GNNs) have emerged as a popular ML tool for learning functions on graphs [Ham20, BBCV21]. GNNs are designed to satisfy permutation invariance (for graph-level output), where the choice of enforcing invariance leads to a tradeoff between efficiency and expressivity. For example, one of the most efficient GNN architectures is the class of message-passing GNNs [KW16, VCC+17, GSR+17], which uses the same function to process each node neighborhood. However, message-passing GNNs can suffer from expressivity issues when learning global properties of graphs. The expressivity of GNNs has been extensively studied using the graph isomorphism test and comparison with the Weisfeiler-Lehman (WL) algorithms [XHLJ18, MRF+19]—a hierarchy of combinatorial graph invariants (See [HV21] for an introduction and [MLM+23] for a comprehensive survey). To mitigate such issues, more powerful GNNs have been proposed, including higher-order GNNs [MRF+19, MBHSL18], subgraph-based GNNs [CCVB20, TZK21, BFL+21, BFZB22, FBBM22], and spectral GNNs [LRZ+22, HVP+22], yet they typically incur higher computation costs due to higher-order tensor operations or additional preprocessing. By using intractable high-order tensors, universality results are established in [MFSL19] for invariant functions and equivariant functions in [KP19]. [^1] Intractability is expected here, as it is unknown whether the graph isomorphism problem can be solved in polynomial time. Beyond the Weisfeiler-Lehman hierarchy, expressivity of GNNs can be studied using equivariant polynomials [PLK+23] or graph spectral invariants [LRZ+22].
@@
Universality results on approximating invariant or equivariant functions on point clouds (with respect to both permutation and Euclidean symmetries) have been established using polynomials [DM20], complete distances [WK23], or a geometric variant of the Weisfeiler-Lehman test [PC22, JBM+23, HAGD23, RKR+23, HAD24]. Moreover, polynomial-time algorithms to achieve universality on point clouds are discussed in these works (as opposed to intractable algorithms for the case of weighted graphs above; this results from the fact that the data is confined to $\mathbb{R}^{d}$ for fixed $d$). More concretely, [HAGD23] proposed an invariant network inspired by the $2$-WL test that separates all point clouds in $\mathbb{R}^{3}$ with $O(n^{4})$ computational complexity, which is further extended to $\mathbb{R}^{d}$ for any $d>1$ with complexity $O(n^{d+1})$ in [HAGD23, RKR+23], and improved in [HAD24] to $O(n^{d})$ complexity (using a different network architecture motivated by [MBHSL19]). Most related to our approach, [HAGD23] also showed that a more efficient architecture inspired by the $1$-WL test, with $O(n^{2})$ computational complexity (for $d=3$), can achieve universality except for an explicitly-described measure-zero set.
@@
Enforcing symmetries in machine learning models. More generally, invariant and equivariant machine learning is a very active research area, where researchers incorporate symmetries into the design of machine learning models [BBCV21, WFVW21, VHY+23]. There are many different approaches, including orbit averaging [EZ21], frame averaging [PABH+21], representation theory [KLT18, GS22, FWFW20], group convolutions [CW16, KT18, CGW19], weight sharing [RSP17], canonicalization [BRW+20, KMZ+23, APS23, Olv23], and invariant theory [BSV22, VHSF+21, GHL21, Had21, PLK+23, VYH+23], among others. The present work considers the invariant theory approach.

### v3

# A Galois theorem for machine learning: Functions on symmetric matrices and point clouds via lightweight invariant features

@@ Abstract
In this work, we present a mathematical formulation for machine learning of (1) functions on symmetric matrices that are invariant with respect to the action of permutations by conjugation, and (2) functions on point clouds that are invariant with respect to rotations, reflections, and permutations of the points. To achieve this, we provide a general construction of generically separating invariant features using ideas inspired by Galois theory. We construct $O(n^{2})$ invariant features derived from generators for the field of rational functions on $n\times n$ symmetric matrices that are invariant under joint permutations of rows and columns. We show that these invariant features can separate all distinct orbits of symmetric matrices except for a measure zero set; such features can be used to universally approximate invariant functions on almost all weighted graphs. For point clouds in a fixed dimension, we prove that the number of invariant features can be reduced, generically without losing expressivity, to $O(n)$, where $n$ is the number of points. We combine these invariant features with DeepSets to learn functions on symmetric matrices and point clouds with varying sizes. We empirically demonstrate the feasibility of our approach on molecule property regression and point cloud distance prediction.
@@ 1 Introduction
Motivated by such considerations of computational scalability, we propose a relaxation of universal approximation. At the cost of throwing out a closed, measure zero set of “bad” point clouds (respectively, weighted graphs), we provide an approach to these learning problems that has a universal approximation guarantee on the remaining “good” generic point clouds (respectively, weighted graphs), and has computational cost bounded by a constant multiple of the size of the input data.
Our method is to look for computationally tractable invariants that contain the same information as a set of generators for the field of invariant rational functions on the data. It is known that such field generators separate orbits generically, that is, away from a Zariski-closed (and thus Euclidean-closed and measure zero) “bad” subset; thus the proposed invariants do as well. The mathematical techniques we employ are general. If $\Gamma$ is a group acting on a topological measure space $X$, and $G$ is a finite index subgroup of $\Gamma$, we provide a recipe to extend a set of invariants that are generically separating for $\Gamma$ to a set of generically separating invariants for $G$. The construction is inspired by the fundamental theorem of Galois theory, but it is not restricted to polynomials. In particular, it applies to classes of functions generally used in machine learning.
A standard argument based on the Stone-Weierstrass theorem then implies that the proposed invariants are universally approximating away from the “bad” subset.
@@
  • We provide a “Galois” theorem for machine learning. If $\Gamma$ is an arbitrary group acting on a topological measure space $X$ and $G$ is a finite index subgroup of $\Gamma$, the theorem shows how to extend a set of invariants that are generically separating for $\Gamma$ to a set of generically separating invariants for $G$. The construction is similar to the fundamental theorem of Galois theory, but it is not restricted to polynomials. In particular, it applies to various classes of functions used in machine learning.
  • Using our Galois theorem, we provide $O(n^{2})$ invariant features that generically separate real symmetric matrices with respect to the action (1) (Theorem 4.4), so they can be used to universally approximate invariant functions on almost all such matrices (Proposition 4.1). The construction hinges on an idea like one that appears in [Thi00] for a slightly different setting.
  • We show that the same features, applied to the Gram matrix of a point cloud, also generically separate (and therefore generically universally approximate invariant functions on) point clouds $V\in\mathbb{R}^{d\times n}$ with respect to the natural $\mathrm{O}(d)\times\mathfrak{S}_{n}$ action (Theorem 6.4 and Proposition 5.1). The result is not surprising, but nevertheless the proof is not trivial and is new to the best of our knowledge.
  • To improve efficiency, we prove that in the point cloud setting we can express an equivalent class of functions using $O(n)$ invariant features for fixed $d$ (Proposition 6.5). Using ideas from low-rank matrix completion, we show that one can use $O(n)$ features to reconstruct the $O(n^{2})$ generically separating invariants from above (Theorem 6.4).
@@ 2 Related work
Invariant theory. Describing the class of functions invariant under a group action is historically the project of invariant theory. Other works to consider, from the point of view of invariant theory, the same or closely related group actions to the ones discussed above (i.e., on weighted graphs or point clouds), include [ACG96, Thi00, KLR22, GHL21, Had21]; the last two of these have applications to ML in mind. These papers compute complete sets of invariants (i.e., generators for the algebra of invariant polynomials), or separating sets (i.e., invariant polynomials whose evaluation can classify orbits to the extent possible) in special cases. However, a complete or even a separating set of invariants for general $n,d$ is expected to be computationally intractable, because such a set would solve the graph isomorphism problem, for which the fastest known algorithm has quasi-polynomial time complexity. The starting point for the present investigation was the recognition that the method of [Thi00, Theorem 11.2]—which gives generators for the field of rational invariants in a closely related situation where the algebra generators remain out of reach—could be used to produce a set of computationally tractable invariants that uniquely identifies a generic (though not a worst-case) symmetric matrix, up to the $\mathfrak{S}_{n}$ conjugation action, and is not meaningfully bigger than the size of the input data. Other works that use field generation as a relaxation of algebra generation to lower a computational hurdle include [BBSK+23, BSGHR24]. More general work on computation of fields of rational invariants includes [MQB99, Kem07, HK07, HL12, KK12, HL16, Kem16].
@@
Machine learning on weighted graphs. Graph neural networks (GNNs) have emerged as a popular ML tool for learning functions on graphs [Ham20, BBCV21]. GNNs are designed to satisfy permutation invariance (for graph-level output), where the choice of enforcing invariance leads to a tradeoff between efficiency and expressivity. For example, one of the most efficient GNN architectures is the class of message-passing GNNs [KW16, VCC+17, GSR+17], which uses the same function to process each node neighborhood. However, message-passing GNNs can suffer from expressivity issues when learning global properties of graphs. The expressivity of GNNs has been extensively studied using the graph isomorphism test and comparison with the Weisfeiler-Lehman (WL) algorithms [XHLJ18, MRF+19]—a hierarchy of combinatorial graph invariants (See [HV21] for an introduction and [MLM+23] for a comprehensive survey). To mitigate the expressivity issues, more powerful GNNs have been proposed, including higher-order GNNs [MRF+19, MBHSL18, PLK+23], subgraph-based GNNs [CCVB20, TZK21, BFL+21, BFZB22, FBBM22], and spectral GNNs [LRZ+22, HVP+22], yet they typically incur higher computation costs due to higher-order tensor operations or additional preprocessing. By using intractable high-order tensors, universality results are established in [MFSL19] for invariant functions and equivariant functions in [KP19]. [^1] Intractability is expected here, as it is unknown whether the graph isomorphism problem can be solved in polynomial time. In contrast, we relax universality to almost-everywhere universality in order to get the number of features to the same order as the size of the data.
@@
Universality results on approximating invariant or equivariant functions on point clouds (with respect to both permutation and Euclidean symmetries) have been established using polynomials [DM20], complete distances [WK23], or a geometric variant of the Weisfeiler-Lehman test [PC22, JBM+23, HAGD23, RKR+23, HAD24]. Moreover, polynomial-time algorithms to achieve universality on point clouds are discussed in these works (as opposed to intractable algorithms for the case of weighted graphs above; this results from the fact that the data is confined to $\mathbb{R}^{d}$ for fixed $d$). More concretely, [HAGD23] proposed an invariant network inspired by the $2$-WL test that separates all point clouds in $\mathbb{R}^{3}$ with $O(n^{4})$ computational complexity, which is further extended to $\mathbb{R}^{d}$ for any $d>1$ with complexity $O(n^{d+1})$ in [HAGD23, RKR+23], and improved in [HAD24] to $O(n^{d})$ complexity (using a different network architecture motivated by [MBHSL19]). Most related to our approach, [HAGD23] also showed that a more efficient architecture inspired by the $1$-WL test, with $O(n^{2})$ computational complexity (for $d=3$), can achieve universality except for an explicitly-described measure-zero set. We reduce the computational complexity to $O(n)$ by applying ideas from low-rank matrix completion.
@@
Enforcing symmetries in machine learning models. More generally, invariant and equivariant machine learning is a very active research area, where researchers incorporate symmetries into the design of machine learning models [BBCV21, WFVW21, VHY+23]. There are many different approaches, including orbit averaging [EZ21], frame averaging [PABH+21], representation theory [KLT18, GS22, FWFW20], group convolutions [CW16, KT18, CGW19], weight sharing [RSP17], canonicalization [BRW+20, KMZ+23, APS23, Olv23], and invariant theory [BSV22, VHSF+21, GHL21, Had21, PLK+23, VYH+23], among others. The present work considers the invariant theory approach. Specifically, we leverage the invariants of a compact Lie group to obtain invariants of a (finite index) subgroup. Our construction is somewhat related to recent works on symmetry breaking [SGM21, LPZK24, XS24], which reduces the symmetries of an equivariant function by augmenting the input to the function with additional symmetry-breaking parameters. Such symmetry-breaking parameters can be learned [SGM21], sampled from a set based on input and output data symmetries [XS24], or random noise (for full symmetry breaking) [LPZK24].
3 A “Galois” theorem for machine learning applications
Let $X$ be a topological measure space with an action by a group $\Gamma$. In our applications, $X$ is a real vector space and $\Gamma$ is a compact Lie group acting orthogonally, although for the purposes of this section, we do not need the action to respect the measure, or even the topological structure; the only purpose these structures serve is to give subsets of $X$ a notion of “closed and measure zero”. Let $f_{1},\dots,f_{r}$ be a set of $\Gamma$-invariant functions from $X$ to some abelian group $\mathbb{F}$ (which in machine learning applications will generally be $\mathbb{R}$ or perhaps $\mathbb{C}$). We say that $f_{1},\dots,f_{r}$ are generically separating for $\Gamma$ if there exists a closed, measure zero, $\Gamma$-invariant subset $B\subset X$ (the “bad set”) such that $x_{1},x_{2}\in X\setminus B$ must lie in the same orbit of $\Gamma$ if $f_{j}(x_{1})=f_{j}(x_{2})$ for all $j$.
Now let $G\subset\Gamma$ be a subgroup of finite index. In this section, we prove an elementary theorem that gives a condition on some $G$-invariant functions $f^{\star}_{1},\dots,f^{\star}_{s}:X\rightarrow\mathbb{F}$ so that if $\Gamma$-invariants $f_{1},\dots,f_{r}$ are generically separating for $\Gamma$, then also $f_{1},\dots,f_{r},f^{\star}_{1},\dots,f^{\star}_{s}$ are generically separating for $G$. We use this theorem in Section 4 to prove that certain small sets of invariants generically separate orbits for the actions of interest in this paper.
The theorem proven here can be seen as an analogue to a certain basic theorem of Galois theory, adapted for machine learning purposes. Indeed, we use Galois theory to prove a parallel result in Section 5, and in an earlier draft of the present work, we also used Galois theory for the Section 4 result. The latter proof is retained in the Appendix for interest.
Suppose $\mathcal{F}$ is a class of functions $X\rightarrow\mathbb{F}$ that is an abelian group under pointwise addition, is closed under the natural action of $\Gamma$ by $(\gamma f)(x):=f(\gamma^{-1}x)$ for $x\in X,\gamma\in\Gamma$, and such that the nonzero elements of $\mathcal{F}$ have closed, measure-zero vanishing sets. In our applications, $\mathcal{F}$ is the class of polynomial functions. But, for example, if $X$ is a $\mathbb{C}$- or $\mathbb{R}$-vector space and $G$ a linear group, then the class of analytic functions also has these properties, and one could also use other classes.222 For readers who may have been puzzled at the fact that we do not impose any assumptions on the way $G$’s action on $X$ interacts with the latter’s topological or measure structure: it is because of the assumptions on $\mathcal{F}$. If $G$’s action on $X$ is very disruptive of the topological and measure structures, then these assumptions on $\mathcal{F}$ are extremely restrictive. For a natural number $s$, denote $[s]:=\{1,\dots,s\}$. The promised “Galois theorem” for machine learning is as follows:
Theorem 3.1.
Let $X,\Gamma,G,\mathcal{F}$ be as above. Suppose $f_{1},\dots,f_{r}:X\rightarrow\mathbb{F}$ are $\Gamma$-invariant functions that are generically separating for $\Gamma$. If $f^{\star}_{1},\dots,f^{\star}_{s}:X\rightarrow\mathbb{F}$ are $G$-invariant functions belonging to $\mathcal{F}$, such that for $\gamma\in\Gamma$ we have
$\gamma f^{\star}_{j}=f^{\star}_{j}\text{ for all }j\in[s]\quad\Rightarrow\quad\gamma\in G,$
then $f_{1},\dots,f_{r},f^{\star}_{1},\dots,f^{\star}_{s}$ are generically separating for $G$.
Note that the hypothesis does not require $f_{1},\dots,f_{r}$ to belong to $\mathcal{F}$, only $f^{\star}_{1},\dots,f^{\star}_{s}$. Below, we will apply this theorem in the special case that $s=1$, so there is a single $f^{\star}$ that is fixed only by the elements of $G$.
Proof.
Because $f^{\star}_{1},\dots,f^{\star}_{s}$ are $G$-invariant, the stabilizer $\Gamma_{j}$ of each $f_{j}$ in $\Gamma$ contains $G$. The hypothesis on the $f^{\star}_{j}$’s can be summarized as
$\bigcap_{j}\Gamma_{j}=G.$
Because, furthermore, $[\Gamma:G]<\infty$, each $\Gamma_{j}$ has finite index in $\Gamma$. Thus, there are only finitely many functions of the form $\gamma f^{\star}_{j}$, $\gamma\in\Gamma$, namely one for each pair $(j,\gamma\Gamma_{j})$ consisting of $j\in[s]$ and a left coset of the stabilizer $\Gamma_{j}$. They all belong to $\mathcal{F}$, because $\mathcal{F}$ is $\Gamma$-stable. Thus, the finitely many functions
$\gamma f^{\star}_{j}-f^{\star}_{j},\;j\in[s],\;\gamma\notin\Gamma_{j}$
all belong to $\mathcal{F}$ as well (because it is an abelian group). Furthermore, they are all nonzero because $\gamma\notin\Gamma_{j}$ for each one. So by the hypothesis on $\mathcal{F}$, they all have closed, measure zero vanishing sets. Let
$B=\bigcup_{j\in[s],\gamma\in\Gamma\setminus\Gamma_{j}}\{x\in X:(\gamma f^{\star}_{j}-f^{\star}_{j})(x)=0\}$ (2)
be the union of these; by the above, this is a union of only finitely many distinct closed, measure zero sets, thus it is closed and measure zero.
Meanwhile, because $f_{1},\dots,f_{r}$ are generically separating for $\Gamma$, we know that there exists another closed, measure zero set $B^{\prime}$ on the complement of which any two distinct $\Gamma$-orbits are separated by some $f_{j}$.
Then $B\cup B^{\prime}$ is closed and measure zero, and we claim that any $x_{1},x_{2}\in X\setminus(B\cup B^{\prime})$ lying in distinct $G$-orbits are separated either by some $f_{j}$ or by some $f^{\star}_{j}$. Indeed, if $x_{1},x_{2}$ lie in distinct $\Gamma$-orbits, then they are separated by some $f_{j}$; while if they lie in the same $\Gamma$-orbit but distinct $G$-orbits, then there exists $\gamma\in\Gamma\setminus G$ with $x_{1}=\gamma x_{2}$. In the latter case, there exists $j\in[s]$ with $\gamma\notin\Gamma_{j}$ because $G=\bigcap_{j}\Gamma_{j}$, and then
$f^{\star}_{j}(x_{1})-f^{\star}_{j}(x_{2})$ $=f^{\star}_{j}(\gamma^{-1}x_{2})-f^{\star}_{j}(x_{2})$
$=(\gamma f^{\star}_{j}-f^{\star}_{j})(x_{2})$
$\neq 0,$
where the final inequality is because $x_{2}\notin B$ (and $B$ contains the vanishing set of $\gamma f^{\star}_{j}-f^{\star}_{j}$ by definition). So $x_{1},x_{2}$ are separated by $f^{\star}_{j}$. ∎
Here is the sense in which this is a Galois theory analogue. Let $\mathbb{F}$ be a field (which we think of as a ground field), and let $\alpha_{1},\dots,\alpha_{r}$ be elements of some (unspecified) extension of $\mathbb{F}$. Let $\mathbb{K}$ be a finite Galois extension of $\mathbb{F}(\alpha_{1},\dots,\alpha_{r})$. This extension has a Galois group, say $\Gamma$, and $\mathbb{F}(\alpha_{1},\dots,\alpha_{r})=\mathbb{K}^{\Gamma}$. Now let $G\subset\Gamma$ be a subgroup of $\Gamma$. Then $\mathbb{K}^{G}$ is a finite extension of $\mathbb{F}(\alpha_{1},\dots,\alpha_{r})$, and we may be interested in identifying generators for this extension. If we can find $\alpha^{\star}_{1},\dots,\alpha^{\star}_{s}\in\mathbb{K}^{G}$ such that
$\gamma\alpha^{\star}_{j}=\alpha^{\star}_{j}\text{ for all }j\in[s]\quad\Rightarrow\quad\gamma\in G,$
then the Fundamental Theorem of Galois Theory tells us that $\alpha^{\star}_{1},\dots,\alpha^{\star}_{s}$ generate $\mathbb{K}^{G}$ as an extension of $\mathbb{F}(\alpha_{1},\dots,\alpha_{r})$, or in other words, that
$\mathbb{F}(\alpha_{1},\dots,\alpha_{r},\alpha^{\star}_{1},\dots,\alpha^{\star}_{s})=\mathbb{K}^{G}.$
In this analogy, the hypothesis that $\alpha_{1},\dots,\alpha_{r}$ generate $\mathbb{K}^{\Gamma}$, and the conclusion that $\alpha_{1},\dots,\alpha_{r},\alpha^{\star}_{1},\dots,\alpha^{\star}_{s}$ generate $\mathbb{K}^{G}$, play the role of the hypothesis that $f_{1},\dots,f_{r}$ are generically separating for $\Gamma$ and the conclusion that $f_{1},\dots,f_{r},f^{\star}_{1},\dots,f^{\star}_{s}$ are generically separating for $G$ in Theorem 3.1 above.
This connection is somewhat more than an analogy, because if $\mathbb{K}$ happens to be the function field of an algebraic variety $X$ over an algebraically closed field $\mathbb{F}$ of characteristic zero, and $\alpha_{1},\dots,\alpha_{r},\alpha^{\star}_{1},\dots,\alpha^{\star}_{s}$ are rational functions on $X$, then Rosenlicht’s Theorem [Ros56, Theorem 2] tells us that they generate the invariant field $\mathbb{K}^{\Gamma}$, respectively $\mathbb{K}^{G}$, if and only if they are generically separating for $\Gamma$, respectively $G$. So Theorem 3.1 is actually a consequence of the Fundamental Theorem of Galois Theory plus Rosenlicht’s theorem in this special case.
But the proof of Theorem 3.1 shows that the conclusion holds for an elementary reason that evades this algebraic machinery, so it can be applied in situations where the functions $f_{j}$ and $f^{\star}_{j}$ are not rational functions, or $X$ is not an algebraic variety; this flexibility is the source of its potential usefulness to machine learning theory. It also gives an explicit description of the “bad set” $B$ introduced by the $f^{\star}_{j}$’s (equation (2)).